DEC Time

07:36
  • Having problems with the AdSec Arbitrator process. It keeps Crashing.
  • Brandon is working on it.
08:12
  • OSCO
  • Center pupils gave us an error.
  • WfsArbitrator _98658|WAR| 86|2022-02-18 08:13:15.868834| WFS > Command returns retryable error : RETRY: Camera lens temperature too low. T= 6.89, minimum allowed is 7.00 (0) NO_ERROR
    WfsArbitrator_98658|INF| 87|2022-02-18 08:13:15.868880| COMMANDHANDLER > Command not executed (status RETRY) (errstr RETRY: Camera lens temperature too low. T= 6.89, minimum allowed is 7.00 (0) NO_ERROR)
    WfsArbitrator_98658|WAR| 88|2022-02-18 08:13:15.868909| MAIN > centerPupils: WARNING - RETRY: Camera lens temperature too low. T= 6.89, minimum allowed is 7.00 (0) NO_ERROR
  • We re-tried center pupils and it worked.
  • F| 139|2022-02-18 08:15:46.277315| WFS > Executing optimizeGain...
    WfsArbitrator_98658|INF| 140|2022-02-18 08:15:46.277369| WFS > Python command: WfsArbScripts.optimizeGain(self, sensor = 'FLAOWFS', instr = 'LUCIFER')
    WfsArbitrator_98658|ERR| 141|2022-02-18 08:15:53.024180| WFS > Command returns error : Data analysis routine (in IDL) did not finish correctly (-3006) VAR_NOT_FOUND_ERROR
    WfsArbitrator_98658|INF| 142|2022-02-18 08:15:53.024233| COMMANDHANDLER > Command execution failed: Fsm forced to Failure
    WfsArbitrator_98658|INF| 143|2022-02-18 08:15:53.024258| WFS > Executing stopLoop...
    WfsArbitrator_98658|INF| 144|2022-02-18 08:15:53.024294| WFS > Python command: WfsArbScripts.stopLoop(self, sensor = 'FLAOWFS', instr = 'LUCIFER')
    WfsArbitrator_98658|INF| 145|2022-02-18 08:15:53.028665| WFS > Command returns success .
    WfsArbitrator_98658|INF| 146|2022-02-18 08:15:53.028692| MASTDIAGN > Setting master diagnostic in open loop mode
    WfsArbitrator_98658|INF| 147|2022-02-18 08:15:53.028805| ADSEC > Executing openLoop...
    WfsArbitrator_98658|INF| 148|2022-02-18 08:15:53.028819| ADSEC > Requesting StopAo to AdSecArbitrator...
    WfsArbitrator_98658|ERR| 149|2022-02-18 08:15:58.946565| MAIN > OptimizeGain: FATAL - Data analysis routine (in IDL) did not finish correctly (-3006) VAR_NOT_FOUND_ERROR
    WfsArbitrator_98658|INF| 150|2022-02-18 08:15:59.814020| COMMANDHANDLER > FSM (status: AOSet) has received command 2023 (EnableDisturb)
    WfsArbitrator_98658|INF| 151|2022-02-18 08:15:59.814159| WFS > Executing EnableDisturb...
    WfsArbitrator_98658|INF| 152|2022-02-18 08:15:59.814246| WFS > Python command: WfsArbScripts.enableDisturb(self, sensor = 'FLAOWFS', instr = 'LUCIFER', wfs = True, ovs = False)
    WfsArbitrator_98658|INF| 153|2022-02-18 08:15:59.815322| WFS > Command returns success .

  • Problem with optimize gain
  • We re-tried.
  • Going now.
  • 20220218_082102
    20220218_082115
    20220218_082128
    20220218_082202
    20220218_082215
    20220218_082226
    20220218_082444
    20220218_082457
    20220218_082510
    20220218_083028
    20220218_083041
    20220218_083053
    20220218_083428
    20220218_083440
    20220218_083452
08:47
  • Heading to a new target.
  • 10 Degrees from the moon.
  • WFS mag is showing 11.5
  • I'm seeing a few rolled edges, I'll try to get some data.
  • 20220218_091001
    20220218_091026
  • Caught you, you bugger.
  • 374 seems to have a detached magnet.
  • Putting in the do not use list for current..
09:32
  • Going again.
  • Major seeing bubble. Optimize gain failed.
09:37
  • AdSec in Failure. This was not a Set Z0.
  • The gains keep getting ramped to ludicrous levels.
  • Sending the preset again.
09:51
  • Seeing exploded, loop opened.
  • We're going to keep LN gain going to do a quick test.
10:00
  • We are doing ESM. We just checked to make sure that the thing wouldn't go to Bin1.
  • It stayed in Bin2, so I think that we are good.
  • 20220218_100922
    20220218_100947
  • Loads 'o bad seeing, we had a couple of safe skips.
  • housekeeper.L |INF| 18340|2022-02-18 10:09:47.520954| SwitchSafeSkipFrameCnt > 10801
    housekeeper.L |INF| 18348|2022-02-18 10:09:48.591405| SwitchSafeSkipFrameCnt > 10971
    housekeeper.L |INF| 18376|2022-02-18 10:09:49.738475| SwitchSafeSkipFrameCnt > 10985
  • I tested the "Force Open Loop" Button. It seemed to work and we can then hit the Re-close Loop button.
  • Stopping for bad seeing.
10:56
  • Seeing has markedly improved.
  • Going to try the bin3 moon test.
  • Reconfiguring.
  • 10 degrees from the moon, we were not correcting anything. Optimize gain kept failing.
  • The gain plots were completely flat.
  • Trying 15 degrees from the moon.
  • This one looks like bin2
  • Network bogging down again.
  • I ssh-ed in to obs 3 and killed my LBTplot and FACSUM. It didn't seem to help.
  • It's stable, but I think there is still an annoying amount of moonlight.
  • Going to see if it ramps up the gain when we resume from a pause.
  • If it ramped, there was no indication.
  • I took a tracking number right when we did a big offset, to see if the gains ramp.
  • 20220218_114753
11:51
  • Making a new script.
12:11
  • Cat's eye nebula
  • Guide star bin2
  • Bad Seeing again!

-- GregoryTaylor - 18 Feb 2022
Topic revision: r1 - 18 Feb 2022, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback